OpenStudy (anonymous):

what are the x and y intercepts of f(x) f(x)=x^(2)-5x+1 x= y=

OpenStudy (bahrom7893):

x-int: set y = 0, solve for x y-int: set x = 0, solve for y

OpenStudy (anonymous):

is this right x=5+sqrt21/2,5-sqrt21/2 y= 1

OpenStudy (bahrom7893):

y is right D = 25-4=21.. yea they're right

OpenStudy (anonymous):

thank you